1.   Find the volume generated under the curve y=x^3 + 3 and the x-axis between x=0 and x=3, rotated about the x-axis.



1
Expert's answer
2020-12-16T19:58:08-0500

The formula of finding the volume of the solid generated by revolving the curve about x axis is found by multiplying π by the integration of the square of the function in given interval.

V=π"\int"03(x3+3)2dx=π"\intop"03(x6+6x3+9)dx=π(x7/7+6x4/4+9x)/03=π(2187/7+486/4+27)=

=π6453/14